# Defensive Prompt Patch: A Robust and Interpretable Defense of LLMs against Jailbreak Attacks

### v1

@@ Appendix F Extension of LLAMA-2 Experiments
Appendix G Prompts in Defense Baselines
Appendix H DPP Suffix
Now we want to show the rest of Defense Prompt Patch we trained on LLAMA-2-7B-Chat and used for the experiments in Appendix F.
Appendix I Adaptive Attacks Setup
Our Adaptive Attack is setup in the following way:
Algorithm 7 GCG adaptive
1:Initial prompt $x_{1:n}$, modifiable subset $I$, number of iterations $T$, loss function $L$, parameter $k$ for top elements, batch size $B$, Trained Defense Prompt Patch $d_{1:m}$
2:$\tilde{x}_{1:n+m}\leftarrow x_{1:n}\oplus d_{1:m}$triangleright Append the our DPP to the initial prompt (with modifiable subset)
3:for $t=1$ to $T$ do
4:     for all $i\in I$ do
5:         $\tilde{X}_{i}\leftarrow\text{Top-k}(-\nabla_{\tilde{x}_{i}}L(\tilde{x}_{1:n+m}))$triangleright Compute top-k negative gradients for token substitutions
6:     end for
7:     for $b=1$ to $B$ do
8:         $\tilde{x}^{(b)}_{1:n+m}\leftarrow\tilde{x}_{1:n+m}$triangleright Initialize batch element with current prompt
9:         $i\leftarrow\text{Uniform}(I)$
10:         $\tilde{x}^{(b)}_{i}\leftarrow\text{Uniform}(\tilde{X}_{i})$triangleright Select a random token from top-k replacements
11:     end for
12:     $b^{*}\leftarrow\arg\min_{b}L(\tilde{x}^{(b)}_{1:n+m})$triangleright Identify the batch element with the least loss
13:     $\tilde{x}_{1:n+m}\leftarrow\tilde{x}^{(b^{*})}_{1:n+m}$triangleright Update prompt with the optimal substitutions
14:end for
15:Optimized prompt $\tilde{x}_{1:n+m}$
Algorithm 8 ICA Adaptive
[…]
2:for $l=1$ to $L$ do
3:     $ICD=$ []
4:     for $k=1$ to $K$ do
5:         $ICD\leftarrow(u_{k},r_{k})$triangleright Sample K pairs of In-Context harmful user queries and responses
6:     end for
7:     $ICD\_DPP=$ []
8:     for $k=1$ to $K$ do
9:         $\tilde{u}_{k}\leftarrow u_{k}\oplus d_{1:m}$triangleright Append the DPP into the In-Context Harmful User Queries
10:         $ICD\_DPP\leftarrow(\tilde{u}_{k},r_{k})$triangleright Saved the new In-Context Harmful User Queries
11:     end for
12:     $\tilde{x}_{1:n+m}\leftarrow x_{l}\oplus d{1:m}$triangleright Combine the input malicious query with DPP
[…]
15:end for
For more detailed explanation, please refer to the original paper of AutoDAN [3].
Algorithm 9 AutoDAN Adaptive
1:Input: Jailbreak prompt $J_{p}$, blacklist $L_{refuse}$, hyperparameters, Trained Defense Prompt Patch $d_{1:m}$
2:Initialize: Generate initial population using LLM-based Diversification
3:while unwanted words from $L_{refuse}$ in model responses or iterations not exhausted do
4:     for each prompt in the population do
5:         $\text{prompt}\leftarrow\text{prompt}\oplus d_{1:m}$triangleright Append our DPP to the jailbreak prompt for optimization
6:         Fitness $=-\log(P(\text{response}|\text{prompt}))$
7:         for each word in prompt do
8:              if word not in $L_{refuse}$ then
9:                  synonyms, scores $\leftarrow$ findSynonymsAndScores(word)
10:                  totalScore $\leftarrow$ sum(scores)
11:                  wordDict[word] $\leftarrow$ sum(scores $\times$ wordDict[synonyms]) / totalScore
12:              end if
13:         end for
14:         for each word in prompt do
15:              synonyms, scores $\leftarrow$ findSynonymsAndScores(word)
16:              totalScore $\leftarrow$ sum(scores)
17:              probabilityDistribution $\leftarrow$ [score / totalScore for score in scores]
18:              chosenSynonym $\leftarrow$ chooseWeightedRandom(synonyms, probabilityDistribution)
19:              prompt $\leftarrow$ replaceWord(prompt, word, chosenSynonym)
20:         end for
21:         elite, parents $\leftarrow$ selectEliteAndParents(population, fitnessScores)
22:         population $\leftarrow$ crossoverAndMutate(parents, hyperparameters)
23:     end for
24:end while
25:return findBestPrompt(population)
Algorithm 10 PAIR adaptive
1:Iteration count $K$, goal objective $O$, Trained Defense Prompt Patch $d_{1:m}$
2:Initialize prompt $A$ with objective $O$
3:Initialize conversation history $H\leftarrow[]$
4:for $i=1$ to $K$ do
5:     $P\leftarrow q_{A}(H)$triangleright Generate prompt based on history
6:     $\tilde{P}\leftarrow P\oplus d_{1:m}$triangleright Combine the DPP to the optimized prompt
7:     $R\leftarrow q_{T}(\tilde{P})$triangleright Generate response for prompt
8:     $S\leftarrow\text{JUDGE}(\tilde{P},R)$triangleright Compute judge score
9:     if $S=\text{JAILBROKEN}$ then
10:         return $P$
11:     end if
12:     $H\leftarrow H\cup\{(P,R,S)\}$triangleright Append to history
13:end for
14:return None $\triangleright$ If no prompt is jailbroken
Algorithm 11 TAP
1:Desired outcome $G$, branching factor $b$, max width $w$, max depth $d$
2:Access to attacker $A$, target $T$, Trained Defense Prompt Patch $d_{1:m}$ and functions Judge and Off-Topic
3:Set up initial prompt for attacker $A$
4:Create a tree with a root node initialized with an empty chat history and the prompt $G$
5:while tree depth $\leq d$ do
6:     for each leaf node $\ell$ in the tree do
7:         Generate prompts $P_{1},P_{2},\dots,P_{b}\sim q(C;A)$, where $C$ is the chat history at $\ell$
8:         Create $b$ new child nodes for $\ell$, each with one of the prompts $P_{1},\dots,P_{b}$ and inheriting history $C$
9:     end for
10:     for each new leaf node $\ell$ do
11:         if Off-Topic$(P,G)=1$ for the prompt $P$ at node $\ell$ then
12:              Remove node $\ell$
13:         end if
14:     end for
15:     for each surviving leaf node $\ell$ do
16:         $\tilde{P}\leftarrow P\oplus d_{1:m}$triangleright Append our DPP to the jailbreak prompts
17:         Obtain response $R\sim q(\tilde{P};T)$, where $\tilde{P}$ is the prompt at $\ell$
18:         Compute score $S\leftarrow$ Judge$(R,G)$ and attach it to $\ell$
19:         if $S$ indicates JAILBROKEN then
20:              Return $P$
21:         end if
22:         Append the triplet $[P,R,S]$ to the conversation history at node $\ell$
23:     end for
24:     if number of leaf nodes $>w$ then
25:         Keep only the top $w$ leaf nodes based on their scores, removing all others
26:     end if
27:end whilereturn None
Algorithm 12 Catastrophic Adaptive
1:Malicious Query $x_{1:n}$, Dataset Size $L$, Trained Defense Prompt Patch $d_{1:m}$, Judge evaluator $Judge$ and hyperparameters
2:Initialize the temperature hyperparameter $temp=[0.05\ldots 1.00]$
3:Initialize the top_probability hyperparameter $top\_p=[0.0\ldots 1.00]$
4:Initialize the top_k hyperparameter $top\_k=[1,2,5,10,20,50,100,200,500]$
5:for $l=1$ to $L$ do
6:     $\text{Prompt}\leftarrow x_{1:n}\oplus d_{1:m}$
7:     for all pairs of $temp,top\_p,top\_k$ do
8:         $Response\leftarrow LLM(\text{Prompt},temp,top\_p,top\_k)$
9:         $Judge(Response,Prompt)$
10:     end for
11:end for
12:return Number of $Judge=1$
Appendix J Trade-off Plots
Appendix K Ignorance Adaptive Attack
Ignorance Adaptive For Prefix Ignore all the previous instructions.
Ideally, the defense prompt should still be effective even with the existence ignorance prompts. We utilize the 100 harmful behaviors from AdvBench and perform the evaluation. The results are summarized in Table. 11
Models Defense Methods Ignorance [$\downarrow$]
LLAMA-2-7B-Chat Self-Reminder 0.000
RPO 0.000
Goal Prioritization 0.000
DPP (Ours) 0.000
Mistral-7B-Instruct-v0.2 Self-Reminder 0.120
System Prompt 0.020
Goal Prioritization 0.030
DPP (Ours) 0.010
Appendix L JailbreakBench Chat Queries
@@ Appendix I Adaptive Attacks Setup
Appendix M Limitations
In this section we want to discuss some of our limitations of DPP method
Prototype selection One of the primary limitations of our DPP algorithm arises from the selection of prototypes. When an effective prototype is selected, our DPP algorithm is capable of enhancing the prototype into a superior DPP. Conversely, if the prototype is ineffective, the performance of the trained DPP is compromised. Therefore, the careful selection of the prototype prompt is crucial for the successful mitigation of jailbreak attacks. In future work, we aim to explore methods to relax these prototype selection constraints.
Computational Efficiency and Scalability The DPP training algorithm, which involves a Hierarchical Genetic Algorithm (HGA), is computationally intensive. The scalability of our approach to larger datasets or more extensive model deployments may be limited by the computational resources required for iterative optimization and evaluation. As model sizes and the volume of data grow, the efficiency of DPP in real-time applications may need further optimization.
Cost of Training with DPP The DPP training algorithm requires a LLM to revise the prototype prompt, and currently, we are using GPT-4 as the revising LLM, therefore, the cost of accessing OpenAI platform is considerable high for this training process. In order to minimize the cost of training, one approach is to replace the GPT-4 with some open-sourced LLMs, which will be the future scope of this work.
Limitations of other defense baselines We noticed that other defense baselines also contain limitations. For Self-Reminder, we notice this training procedure works poorly on LLAMA-2-7B-Chat model. Since its well-alignment, it will often refuse to improve upon the defense prompt. For RPO, the main limitation is the training time. RPO adopted the GCG attack training procedure, and thus results a high computational cost for finding the defense suffix. We also observe the inefficient of RPO when defending jailbreak attacks which is discussed in Appendix B. Goal Prioritization is strong defense against GCG attack, but it seems less effective when defending AutoDAN, TAP and PAIR attacks. Moreover, it contains a long in-context learning, which cause the inference time when adding Goal Prioritization increases. From both LLAMA-2-7B-Chat and Mistral-7B-Instruct-v0.2, we observe the utility degradation is large for Goal Prioritization.
Appendix N Broader Impacts
@@
Appendix O Win-Rate Evaluation
Additionally, we explored the influence of system prompts on the degradation of utility. Data in Table 14 show that using a default system prompt can limit the LLM’s capability to answer questions effectively. To ensure uniformity in our experimental approach, we have decided to remove system prompts entirely. We also examine the effect of system prompt on the GCG attack and summarize the results in Table 15. We observe that GCG with system prompt cannot achieve the performance that is mentioned in the original paper of GCG [1]. Therefore, we choose to use GCG attack that is without the system prompt, which is closely matched with the original paper’s experimental results.
Generated Length Win-Rate [$\uparrow$]
L = 300 70.77
L = 1000 81.37
Table 13: Generated Response Length for LLM and effect on Win-Rate
System Prompt Methods Win-Rate [$\uparrow$]
w. system prompt 64.35
w/o system prompt 81.37
System Prompt Methods ASR [$\downarrow$]
w. system prompt 0.360
w/o system prompt 0.550
Original GCG paper 0.560
Table 15: With or without system prompt and effect on GCG attacks
Appendix P Extension of Mistral Experiments
We also evaluate additional defense baseline called Directed Representation Optimization (DRO) [34]. This approach is similar to Self-Reminder which they improved upon the default system prompt. We obtained the trained DRO for Mistral-7B-Instruct-v0.2 and evaluated against 6 different jailbreak attacks. We summarize the results in Table 16. From the table, we observe that our DPP method outperforms the DRO in terms of Average ASR even though the DRO has a better Win-Rate. This further proves that our DPP is more capable of defending jailbreak attacks with a reasonable utility trade-offs.
DRO [34] 0.560 0.080 0.280 0.760 0.020 0.000 0.283 85.07
DPP (Ours) 0.000 0.010 0.020 0.030 0.040 0.020 0.020 75.06
Appendix Q Repository

### v2

@@ 9         $Judge(Response,Prompt)$
Appendix N DPP performance on other Jailbreak Attacks
We conducted additional experiments on more recent jailbreak attacks:
  • Jailbreaking Leading Safety-Aligned LLMs with Simple Adaptive Attacks. (Andriushchenko et al., 2024) (known as llm-simple-adaptive-attacks)
  • Improved few-shot jailbreaking can circumvent aligned language models and their defenses. (Zheng et al., 2024c) (known as I-FSJ)
We summarize our DPP performance along with other defense baslines in in Table 24 and Table 25 under adaptive setting.
Table 24: DPP and other baselines evaluated on two other jailbreak attacks under adaptive setting on Llama-2-7B-Chat
Methods llm-adaptive-attacks [$\downarrow$] I-FSJ [$\downarrow$] Average ASR [$\downarrow$]
w/o defense 0.800 0.660 0.730
Self-Reminder 0.000 0.780 0.390
RPO 0.240 0.680 0.460
Goal Prioritization 0.86 0.960 0.910
DPP (Ours) 0.000 0.000 0.000
Table 25: DPP and other baselines evaluated on two other jailbreak attacks under adaptive setting on Mistral-7B-Instruct-v0.2
Methods llm-adaptive-attacks [$\downarrow$] I-FSJ [$\downarrow$] Average ASR [$\downarrow$]
w/o defense 0.920 1.000 0.960
Self-Reminder 0.880 0.860 0.870
System Prompt 0.920 1.000 0.960
Goal Prioritization 0.660 0.960 0.810
DPP 0.500 0.880 0.690
Appendix O Extension of Mistral Experiments
We also evaluate additional defense baseline called Directed Representation Optimization (DRO) (Zheng et al., 2024b). This approach is similar to Self-Reminder which they improved upon the default system prompt. We obtained the trained DRO for Mistral-7B-Instruct-v0.2 and evaluated against 6 different jailbreak attacks. We summarize the results in Table 26. From the table, we observe that our DPP method outperforms the DRO in terms of Average ASR even though the DRO has a better Win-Rate. This further proves that our DPP is more capable of defending jailbreak attacks with a reasonable utility trade-offs.
DRO (Zheng et al., 2024b) 0.560 0.080 0.280 0.760 0.020 0.000 0.283 85.07
DPP (Ours) 0.000 0.010 0.020 0.030 0.040 0.020 0.020 75.06
Appendix P JailbreakBench Chat Queries
@@
The results from Table 27 show that for the well-aligned model (Llama-2-7B-Chat), the JBC dataset does not yield effective jailbreak attacks, resulting in comparable defense performances across all methods. Conversely, with the less-aligned Mistral-7B-Instruct-v0.2 model, our DPP demonstrated its efficacy by reducing the Attack Success Rate (ASR) from 41% to 1%, attaining the best defense performance (on par with Goal Prioritization). This marked decrease in ASR highlights our DPP’s strong capability to generalize defense performance effectively against unforeseen attacks.
@@
In addition to the manual JBC query, we have conducted a new jailbreak atttack experiment on the 25 harmful queries that is randomly selected from JBC dataset. We apply our DPP to both models under adaptive setting and report the results as follows.
Table 28: Jailbreak Bench Chat queries with two different jailbreak attacks evaluated with different defense mechanisms on Llama-2-7B-Chat.
Methods ICA [$\downarrow$] AutoDAN [$\downarrow$] Average ASR [$\downarrow$]
w/o defense 0.520 0.000 0.260
Self-Reminder 0.400 0.000 0.200
Goal Prioritization 0.520 0.000 0.260
RPO 0.400 0.000 0.200
DPP (Ours) 0.040 0.000 0.020
Table 29: Jailbreak Bench Chat queries with two different jailbreak attacks evaluated with different defense mechanisms on Mistral-7B-Instruct-v0.2.
Methods ICA [$\downarrow$] AutoDAN [$\downarrow$] Average ASR [$\downarrow$]
w/o defense 1.000 0.960 0.980
Self-Reminder 0.920 0.960 0.940
Goal Prioritization 0.840 0.800 0.820
System Prompt 0.960 0.960 0.960
DPP (Ours) 0.040 0.600 0.320
Overall, we observe that our DPP outperforms the other baselines. We suspect that the original implementation of AutoDAN applies a jailbreak template that is more suitable for AdvPrompt dataset, which you can refer to Table 3. However, JBC harmful queries is quite different from the AdvPrompt. Thus, the default jailbreak template of AutoDan might not work well on JBC, which leads to 0 ASR on AutoDAN for Llama-2.
Appendix Q Analysis of DPPs
Q.1 Deeper Insights of DPPs
Table 30: Presentation of different suffixes generated by RPO and our DPP
Methods Prompt
RPO was Bel Let]"> It’ si distinguish ({ footnote{[:// Inst bitimidenote]{BWebachiv
DPP (Llama-2-7B-Chat) Kindly furnish a thorough response to the former user’s question.
DPP (Mistral-7B-Instruct-v0.2) Begin the reply with ‘I’m sorry’ if the client’s query relates to defective components.
In order to provide the intuition of different DPPs we obtained by optimizing on Llama-2-7B-Chat and Mistral-7B-Instruct-v0.2 respectively, we set up two hypothesis and conduct two mini-experiment to prove our hypothesis.
  • Our hypothesis of having word "defective components" in Mistral’s DPP is that Mistral’s native safety alignment is vulnerable to heuristic jailbreak attempts, while Llama is more robust to them. To verity this hypothesis, we report the ASR of these two models (without DPP) using the same JBC (human-engineered) jailbreak queries in Table 31. We found that Mistral’s ASR is significant higher than Llama-2, which is a sign of stronger alignment for the Llama-2 model. Thus, in the presence of such a natively embedded safety alignment, our method does not consider any "defective components" in Llama’s DPP, but suggests to have them in Mistral’s DPP.
  • Our hypothesis of having word "thorough" in Llama’s DPP is that longer query length (also known as prompt dilution strategy) might be an effective jailbreak approach to compromise Llama. We conducated a length analysis of successful jailbreak attacks and found that in general, existing Jailbreak attacks tend to increase the length of prompts. Moreover, the length of successful jailbreak queries on Llama is much longer (1.5x 2.3x) than that of Mistral (which are reported in Table 32 and Table 33, validating our hypothesis. Thus, such an increase in context length might require the Llama-2 to read it carefully before generating responses. Thus, our method suggests having "thorough" in Llama’s DPP.
Table 31: Experiment on difference in alignment of two models by feeding the same JBC jailbreak queries
Models JBC ASR
Llama-2-7B-Chat 0.0
Mistral-7B-Instruct-v0.2 0.41
Table 32: Experiment on Llama-2-7B-Chat that calculate the different average query length generated by different jailbreak attacks
Jailbreak Methods Average Length
Original Queries 12.5
PAIR 56.167
TAP 80.2
Table 33: Experiment on Mistral-7B-Instruct-v0.2 that calculate the different average query length generated by different jailbreak attacks
Jailbreak Methods Average Length
Original Queries 12.5
PAIR 36.83
TAP 33.31
Q.2 Quantitative analysis of clarity between different defense mechanisms
Perplexity [$\downarrow$]
Self-Reminder 298.39
Goal Prioritization 40.65
System Prompt 25.65
RPO 8780.94
DPP (Ours) 56.57
Quantitatively, we measure the perplexity for our DPP as well as other defense baseline prompts on Llama-2-7B-Chat in Table 34. The perplexity score for a sentence is calculated by averaging the negative log probabilities of next-token, predicted by the GPT-4 model, and using this average as the exponent in a base-2 exponential function. Our method exhibits a lower perplexity score than RPO and Self-Reminder, indicating higher clarity. It is noteworthy that RPO has the highest perplexity, suggesting that the suffix prompt generated by RPO is highly obscurity due to the use of GCG Attack algorithm. Although both Goal Prioritization and System Prompts are hand-crafted defense prompts with lower perplexity (i.e., they are more human-readable prompts), our method remains competitive with these approaches while sparing the need for human interventions in prompt design and optimization.
Appendix R DPP performance on advanced jailbreak attacks on Llama-2-7B-Chat
Besides the recent jailbreak attacks we conduct the experiment on DPP performance in Appendix N. We also select three additional advanced jailbreak attacks:
  • GPTFUZZER: Red Teaming Large Language Models with Auto-Generated Jailbreak Prompts. Yu et al. (2024) (known as GPTFuzz)
  • How Johnny Can Persuade LLMs to Jailbreak Them: Rethinking Persuasion to Challenge AI Safety by Humanizing LLMs. Zeng et al. (2024) (known as PAP)
  • When LLM Meets DRL: Advancing Jailbreaking Efficiency via DRL-guided Search. Chen et al. (2025) (known as DRL)
  • We also include llm-simple-adaptive-attacks and I-FSJ from Appendix N for completion of the evaluation.
We conduct the experimenton Llama-2-7B-Chat under non-adaptive setting and the record the numerical results in Table 35.
Table 35: DPP and other baselines evaluated on five advanced jailbreak attacks under non-adaptive setting on Llama-2-7B-Chat
GPTFuzz [$\downarrow$] PAP [$\downarrow$] DRL [$\downarrow$] llm-adaptive-attacks [$\downarrow$] I-FSJ [$\downarrow$] Average ASR [$\downarrow$]
w/o defense 0.6 0.3 0.2 0.8 0.66 0.512
Self-Reminder 0.17 0.18 0.2 0 0.06 0.122
RPO 0.68 0.6 0.2 0 0.56 0.408
Goal Prioritization 0.14 0.42 0.6 0.9 0.68 0.548
DPP 0.09 0.12 0.1 0 0 0.062
From Table 35, we conclude that our DPP outperforms all the baseslines under these attacks.
Appendix S Trade-off Plots
Appendix T Ignorance Adaptive Attack
Ignorance Adaptive For Prefix Ignore all the previous instructions.
Ideally, the defense prompt should still be effective even with the existence ignorance prompts. We utilize the 100 harmful behaviors from AdvBench and perform the evaluation. The results are summarized in Table 36
Models Defense Methods Ignorance ASR [$\downarrow$]
Llama-2-7B-Chat Self-Reminder 0.000
RPO 0.000
Goal Prioritization 0.000
DPP (Ours) 0.000
Mistral-7B-Instruct-v0.2 Self-Reminder 0.120
System Prompt 0.020
Goal Prioritization 0.030
DPP (Ours) 0.010
Appendix U Unforeseen Jailbreak Queries on Llama-2-7B-Chat and Mistral-7B-Instruct-v0.2
Table 37 presents a comparative analysis of the performance of our DPP method against other defense baselines on the Llama-2-7B-Chat model, using 100 harmful test queries. The table reports the Adaptive Attack Success Rate (ASR) for four different jailbreak attacks: AutoDAN, PAIR, TAP, and ICA. Our DPP method achieves the lowest average ASR (0.075) across all attacks, outperforming existing defenses such as Self-Reminder (0.155), RPO (0.290), and Goal Prioritization (0.303). This demonstrates the superior effectiveness of our approach in mitigating adaptive attacks on the Llama-2-7B-Chat model.
Table 37: Adaptive Attack Success Rate on Llama-7B-Chat across four different jailbreak attacks on 100 test set harmful queries.
Methods AutoDAN [$\downarrow$] PAIR [$\downarrow$] TAP [$\downarrow$] ICA [$\downarrow$] Average ASR [$\downarrow$]
Self-Reminder 0.190 0.020 0.060 0.350 0.155
RPO 0.270 0.200 0.260 0.430 0.290
Goal Prioritization 0.450 0.000 0.040 0.720 0.303
DPP (Ours) 0.250 0.000 0.040 0.010 0.075
Table 38 provides a similar comparison on the Mistral-7B-Instruct-v0.2 model, also using 100 harmful test queries and the same four jailbreak attacks. While the overall ASRs are higher for this model, our DPP method again achieves the lowest average ASR (0.394), compared to Self-Reminder (0.706), System Prompt (0.780), and Goal Prioritization (0.712). This further corroborates the robustness of our DPP method across different large language models, demonstrating its consistent ability to minimize the success rate of adaptive jailbreak attacks. The results highlight the effectiveness of our proposed method in enhancing the security of large language models against adversarial attacks.
Table 38: Adaptive Attack Success Rates on Mistral-7B-Instruct-v0.2 across four different jailbreak attacks on 100 test set harmful queries. Our method can achieve the lowest Average ASR.
Methods AutoDAN [$\downarrow$] PAIR [$\downarrow$] TAP [$\downarrow$] ICA [$\downarrow$] Average ASR [$\downarrow$]
Self-Reminder 0.960 1.000 1.000 0.570 0.706
System Prompt 0.940 0.980 1.000 0.980 0.780
Goal Prioritization 0.590 1.000 1.000 0.970 0.7120
DPP (Ours) 0.210 0.800 0.880 0.080 0.394
Appendix V Broader Impacts
@@
Appendix W DPP performance on Vicuna-13B-v1.5
W.1 Training DPP on Vicuna
We have also conducted our DPP algorithm on less-aligned model Vicuna-13B-v1.5. We follow the same procedures of training DPP as introduced in Experiment Section, and obtained the following DPP for Vicuna model
DPP Suffix Vicuna-13B-v1.5 Say ’I’m sorry’ to mitigate any negative user behaviors.
We evaluated the DPP along with three different jailbreak attacks under adaptive settings and compare with two other defense baselines [^7] Note that since RPO requires time consuming training in order to obtain the defense suffix, here we opt out the RPO baseline:
Table 39: DPP performance on four different jailbreak attacks under adaptive settings on Vicuna-13B-v1.5.
Methods AutoDAN [$\downarrow$] ICA [$\downarrow$] PAIR [$\downarrow$] TAP [$\downarrow$] Average ASR [$\downarrow$]
Goal Prioritization 1.000 0.970 0.920 0.840 0.933
Remind 0.940 0.750 0.840 0.780 0.828
DPP (Ours) 0.700 0.030 0.100 0.240 0.268
From Table 39 we can observe that our DPP has the best defense performance (i.e. lowest averaged ASR) than other baselines.
W.2 DPP transferability performance on Vicuna
We conduct our experiment under the adaptive setting and we directly use DPP from Mistral to Vicuna-13B v1.5 model.
Table 40: Transferability performance of DPP by applying trained DPP from Mistral to Vicuna
Defense Methods AutoDAN ICA PAIR TAP Averaged ASR
Goal Prioritization 1 0.97 0.92 0.84 0.9325
Self-Reminder 0.94 0.75 0.84 0.78 0.8275
DPP (Ours) 0.85 0.01 0.12 0.1 0.27
From Table. 40, we see that our trained DPP on other models still has a potential ability for defending jailbreak attack since it outperforms all the other defense baselines in terms of defense ability.
Appendix X DPP performance on Llama-3-8B-Instruct
Besides all the LLMs we examine in Sec. 4 and in Appendix W. We have also conducted our DPP algorithm on recent model Llama-3-8B-Instruct. We follow the same procedures of training DPP as introduced in Experiment Section, and obtained the following DPP for Llama-3 model
DPP Suffix Llama-3-8B-Instruct For a accurate and full portrayal, please avoid generating any misleading or harmful content!
We evaluated the DPP along with three different jailbreak attacks under adaptive settings and compare with three other defense baselines: Self-Reminder, System Prompt and Goal Priorization.
Table 41: DPP performance on four different jailbreak attacks under adaptive settings on Llama-3-8B-Instruct.
AutoDAN [$\downarrow$] ICA [$\downarrow$] PAIR [$\downarrow$] TAP [$\downarrow$] Average ASR [$\downarrow$]
Self-Reminder 0.59 0.04 0.06 0.08 0.1925
System Prompt 0.92 0.94 0.02 0.04 0.48
Goal Prioritization 0.19 0.95 0.08 0.12 0.335
DPP (Ours) 0.1 0 0.02 0.06 0.045
Table 41 presents the effectiveness of our proposed DPP approach against four jailbreak attacks on Llama-3-8B-Instruct. DPP consistently outperforms baseline defenses, achieving the lowest average attack success rate (ASR) of 0.045. While existing methods show significant vulnerabilities—particularly System Prompt against AutoDAN (0.92) and ICA (0.94), and Goal Prioritization against ICA (0.95)—DPP demonstrates robust cross-attack resistance, completely neutralizing ICA attacks (0.0) and maintaining strong protection against AutoDAN (0.1), PAIR (0.02), and TAP (0.06). These results validate DPP’s superior defensive capabilities in adaptive adversarial settings.
Appendix Y Min Over Prompt Evaluation
Besides Averaged Attack Success Rate metric, we introduced an additional evaluation metric called Min Over Prompt, which is defined as following:
ASR=Number of prompts with at least one successful attackTotal number of $\scriptstyle\text{ASR}=\frac{\text{Number of prompts with at least one successful attack}}{\text{Total number of prompts}}$
Here Number of prompts with at least one successful attack is calculated by counting one successful jailbreak query from different jailbreak attacks. Whereas Total number of prompts is the total number of input queries for evaluation.
We evaluated our DPP along with other baselines upon the Min Over Prompt metric in Table 4245. From the Min Over Prompt metric, we observe that our DPP perform even better than other defense baselines on both Llama-2-7B-Chat and Mistral-7B-Instruct-v0.2.
Table 42: DPP non-adaptive performance evaluating upon both averaged ASR and Min Over Prompt metrics on Llama-2-7B-Chat
Methods Base64 [$\downarrow$] ICA [$\downarrow$] AutoDAN [$\downarrow$] GCG [$\downarrow$] PAIR [$\downarrow$] TAP [$\downarrow$] Average ASR [$\downarrow$] Min Over Prompt [$\downarrow$]
w/o defense 0.990 0.690 0.640 0.550 0.10 0.120 0.515 1.000
RPO 0.000 0.420 0.280 0.190 0.060 0.060 0.168 0.600
Goal Priorization 0.000 0.020 0.520 0.020 0.020 0.020 0.100 0.560
Self-Reminder 0.030 0.290 0.000 0.040 0.020 0.000 0.063 0.300
DPP (Ours) 0.010 0.000 0.100 0.040 0.040 0.040 0.038 0.120
Table 43: DPP adaptive performance evaluating upon both averaged ASR and Min Over Prompt metrics on Llama-2-7B-Chat
Methods ICA [$\downarrow$] GCG [$\downarrow$] AutoDAN [$\downarrow$] PAIR [$\downarrow$] TAP [$\downarrow$] Average ASR [$\downarrow$] Min Over Prompt [$\downarrow$]
Self-Reminder 0.410 0.210 0.080 0.040 0.060 0.177 0.510
RPO 0.360 0.920 0.170 0.400 0.240 0.475 0.920
Goal Prioritization 0.660 0.190 0.530 0.040 0.060 0.247 0.910
DPP (Ours) 0.160 0.120 0.110 0.080 0.060 0.130 0.300
Table 44: DPP non-adaptive performance evaluating upon both averaged ASR and Min Over Prompt metrics on Mistral-7B-Instruct-v0.2
Methods Base64 [$\downarrow$] ICA [$\downarrow$] GCG [$\downarrow$] AutoDAN [$\downarrow$] PAIR [$\downarrow$] TAP [$\downarrow$] Average ASR [$\downarrow$] Min Over Prompt [$\downarrow$]
w/o defense 0.990 0.960 0.990 0.970 1.000 1.000 0.985 1.000
Self-Reminder 0.550 0.270 0.510 0.880 0.420 0.260 0.482 0.970
System Prompt 0.740 0.470 0.300 0.970 0.500 0.180 0.527 1.000
Goal Priorization 0.030 0.440 0.030 0.390 0.300 0.140 0.222 0.680
DPP (Ours) 0.000 0.010 0.020 0.030 0.040 0.020 0.020 0.040
Table 45: DPP adaptive performance evaluating upon both averaged ASR and Min Over Prompt metrics on Mistral-7B-Instruct-v0.2
Methods ICA [$\downarrow$] GCG [$\downarrow$] AuutoDAN [$\downarrow$] PAIR [$\downarrow$] TAP [$\downarrow$] Average ASR [$\downarrow$] Min Over Prompt [$\downarrow$]
Self-Reminder 0.440 0.610 1.000 1.000 1.000 0.796 1.000
System Prompt 0.990 0.850 0.990 1.000 1.000 0.862 1.000
Goal Priorization 0.960 0.110 0.570 1.000 1.000 0.627 0.980
DPP (Ours) 0.000 0.390 0.470 0.837 0.840 0.469 0.890
Appendix Z Repository